# Born-Infeld Electrogravity and Dyonic Black Holes

### v1

@@ V.4 The smoothing of singularities
As mentioned previously, the choice of an imaginary $\beta$ in the determinantal Lagrangian turns the picture $g$-electrodynamics into its standard Born-Infeld form. In fact, the electric field of the point charge becomes regular at the origin; thus, the field invariants $S$ and $P$ will not diverge at the origin as long as the magnetic monopole $p$ is zero. Then, it is natural to wonder about the impact of this regularity on the geometry of spacetime. First, note that the condition $p=0$ has no impact on the geometry because $p$ enters the geometry in the combination $(q^{2}+p^{2})^{1/2}$, since that is the way $q$ and $p$ appear in the energy-momentum tensor (65) as a factor in $r_{o}$. Second, the regular behavior of the electromagnetic field does not prevent the geometry from diverging. A direct inspection of Eq. (66) shows that the curvature scalar $R$ diverges regardless of whether $\beta$ is real or imaginary in the determinantal Lagrangian. If $\beta$ is real (anomalous case), then $R$ diverges at $\rho=1$ ($r=r_{o}$). If $\beta$ is imaginary (standard case), then $R$ is obtained from Eq. (66) by replacing $\beta^{2}$ with $\beta^{2}$ (that is, $r_{o}^{4}\rightarrow-r_{o}^{4}$) and $\epsilon$ with $\epsilon$; thus, $R$ becomes regular at $\rho=1$ but still diverges at $\rho=0$.
@@
Therefore, Born-Infeld electrogravity, both in its anomalous and standard versions, alleviates the singular behavior of the geometry at $\rho=1$ and $\rho=0$ respectively. We can confirm this property at the level of the Riemann tensor by means of the Kretschmann scalar $K=R_{\alpha\beta\gamma\delta}R^{\alpha\beta\gamma\delta}$. Its expression in terms of $f(r)$ and its derivatives is given by
$K=\frac{[1-f(\rho)]^{2}+\rho^{2}\,f^{\prime}(\rho)^{2}+\rho^{4}\,f^{\prime\prime}(\rho)^{2}}{r_{o}^{4}\ \rho^{4}}\ .$ (88)
@@
$K=\frac{1}{\epsilon^{2}\,(\rho-1)}-\frac{4}{3}\,\frac{1+3\,\alpha\,C-2\,\Phi(1)}{\epsilon^{2}\,\sqrt{\rho-1}}+\mathcal{O}(1)\ ;$ (89)
@@
$K=\frac{4}{3}\,\frac{(3\,\alpha\,C-2\,\Psi(0))^{2}}{\epsilon^{2}\,\rho^{6}}+\mathcal{O}(\rho^{-5})\ .$ (90)
@@
The divergent behavior of $K$ in the standard case of Eq. (90) is also better than in the Reissner-Nordström geometry, since $K$ keeps the behavior $\rho^{-6}$ characteristic of the (chargeless) Schwarzschild geometry. Noticeably, this behavior can still be improved by choosing a mass such that  [^9] This mass is extremal for $\alpha=1$ (see Eq. (83)).
$C=\frac{2\ \Psi(0)}{3\ \alpha}\ .$ (92)
@@
$K=\frac{1}{\epsilon^{2}\,\rho^{4}}-\frac{2}{3\epsilon^{2}\,\rho^{2}}+\frac{6}{5\epsilon^{2}}+\mathcal{O}(\rho^{2})\ .$ (94)
Instead, the behavior of $K$ in the anomalous case cannot be improved by choosing a particular value for $C$, since the first term in Eq. (89) does not depend on $C$. Neither the behavior of $K$ in Reissner-Nordström geometry could be improved in this way. Consistently, the same $r^{-4}$ dependence of the Kretschmann scalar was found in [28] solving the Einstein equations sourced by the Born-Infeld electromagnetism, for a dyonic black hole and a critical value of the mass.
Finally, the existence of a true singularity is confirmed by the fact that both $\rho=1$ in the anomalous case, or $\rho=0$ in the standard case can be reached in a finite proper time. This is because the behavior of $f$ at the singularity is either regular (anomalous case) or the same as in an ordinary black hole (standard case). Therefore, spacetime is geodesically incomplete.
@@ VI Conclusions
We have analyzed a (determinantal) Born-Infeld-like theory in which the gravitational and electromagnetic sectors are coupled through a single determinant. The field equations were obtained by independent variations of the metric, the affine connection and the electromagnetic potential using the Palatini formalism. The metric variation gives a constraint equation (reflecting the absence of metric derivatives in the Lagrangian), while the dynamical content comes from the variations with respect to the connection and the potential. Working with the symmetric part of the Ricci tensor allowed us to set the torsion to zero, then the connection equation enforces metric compatibility and fixes the connection uniquely to the Levi–Civita one.
@@
The field equations are naturally written in terms of the inverse of an auxiliary tensor $\bar{q}^{\mu\nu}$, so to express the dynamics in terms of the physical fields, we considered two complementary formulations. On the one hand, the theory reads more naturally as standard Born-Infeld electrodynamics on an effective background defined by the symmetric tensor $\mathcal{G}_{\mu\nu}=g_{\mu\nu}+\epsilon R_{(\mu\nu)}$. On the other hand, by combining the dynamical equations, one obtains an equivalent formulation in the physical metric $g_{\mu\nu}$; this second description displays a sign change in the electromagnetic invariants and therefore we referred to it as “anomalous”. We showed that the two pictures, $\mathcal{G}$” and “gg”, are related by a set of identities that map the pseudoscalars built with $\mathcal{G}$ into those built with $g$ and that also relate the corresponding volume elements (Eqs. (47) and (49)). In particular, these relations explain how the apparent coupling of the electromagnetic field to curvature in picture $\mathcal{G}$ (Eq. (37)) is reabsorbed and translated into a minimally coupled Born-Infeld structure in picture $g$, in agreement with the equivalence principle.
@@
To explore the implications of the model on the singular behavior of the solutions, we analyzed spherically symmetric dyonic configurations. The electrodynamics in the anomalous case (real $\beta$) leads to an electric field that becomes singular at a finite radius $r_{o}$, restricting the domain of the solution to $r>r_{o}$, while the magnetic field keeps its Coulomb form. If instead the parameter $\beta$ is taken imaginary (which is allowed because only even powers of $\beta$ appear in the Lagrangian), the electric field takes the standard Born-Infeld form but the magnetic monopole is divergent at the origin. Remarkably, even though the electric and magnetic monopoles enter the field invariants $S$ and $P$ unequally, they contribute to the energy and pressure on an equal footing through the combination $\sqrt{q^{2}+p^{2}}$, a feature that is inherited by the geometry. The obtained geometric solution approaches the Reissner–Nordström form asymptotically (with an effective coupling replacing Newton’s constant), and the family of solutions admits two, one (extremal) or no horizons depending on the mass and parameter choices.
@@
Concerning singularities, Born-Infeld electrogravity alleviates the geometric divergences compared to the Reissner–Nordström case, but nevertheless spacetime remains geodesically incomplete. The improvement is more pronounced in the (real $\beta$) anomalous case, in which the curvature singularity is shifted from the center to a sphere at $r=r_{o}$, mitigating the pathological inverse power behavior of the Kretschmann scalar. Notably, for imaginary $\beta$ there exists a solution for which the metric becomes regular at the origin and the singularity in the Kretschmann scalar is of order $r^{-4}$.
@@ Appendix B The projective mode
The antisymmetry in the last two indices of $R^{a}{}_{bcd}$ reflects the 2-form nature of $\widetilde{\widetilde{\mathcal{R}}}{}^{a}_{\;b}$, while the indices $a$, $b$ label the different members of the family.
Let us now consider a projective transformation of the connection

### v2

@@ V.5 The smoothing of singularities
As mentioned previously, the choice of an imaginary $\beta$ in the determinantal Lagrangian turns the picture $g$-electrodynamics into its standard BI form. In fact, the electric field of the point charge becomes regular at the origin; thus, the field invariants $S$ and $P$ will not diverge at the origin as long as the magnetic monopole $p$ is zero. Then, it is natural to wonder about the impact of this regularity on the geometry of spacetime. First, note that the condition $p=0$ has no impact on geometry because $p$ enters the geometry in the combination $(q^{2}+p^{2})^{1/2}$, since that is the way $q$ and $p$ appear in the energy-momentum tensor (65) as a factor in $r_{o}$. Second, the regular behavior of the electromagnetic field does not prevent the geometry from diverging. A direct inspection of Eq. (66) shows that the curvature scalar $R$ diverges regardless of whether $\beta$ is real or imaginary in the determinantal Lagrangian. If $\beta$ is real (anomalous case), then $R$ diverges at $\rho=1$ ($r=r_{o}$). If $\beta$ is imaginary (standard case), then $R$ is obtained from Eq. (66) by replacing $\beta^{2}$ with $\beta^{2}$ (that is, $r_{o}^{4}\rightarrow-r_{o}^{4}$) and $\epsilon$ with $\epsilon$; thus, $R$ becomes regular at $\rho=1$ but still diverges at $\rho=0$.
@@
Therefore, BI electrogravity, both in its anomalous and standard versions, alleviates the singular behavior of the geometry at $\rho=1$ and $\rho=0$ respectively. We can confirm this property at the level of the Riemann tensor by means of the Kretschmann scalar $K=R_{\alpha\beta\gamma\delta}R^{\alpha\beta\gamma\delta}$. Its expression in terms of $f(\rho)$ and its derivatives is given by
$K=\frac{4\,[1-f(\rho)]^{2}+4\,\rho^{2}\,f^{\prime}(\rho)^{2}+\rho^{4}\,f^{\prime\prime}(\rho)^{2}}{r_{o}^{4}\ \rho^{4}}\ .$ (95)
@@
$K=\frac{1}{\epsilon^{2}\,(\rho-1)}-\frac{4}{3}\,\frac{1+3\,\alpha\,C-2\sqrt{\pi}\ \frac{\Gamma(\frac{5}{4})}{\Gamma(\frac{3}{4})}}{\epsilon^{2}\,\sqrt{\rho-1}}+\mathcal{O}(1)\ ;$ (96)
@@
$K=\frac{4}{3\pi}\,\frac{\left(3\sqrt{\pi}\,\alpha\,C-2\,\Gamma(\frac{1}{4})\Gamma(\frac{5}{4})\right)^{2}}{\epsilon^{2}\,\rho^{6}}+\mathcal{O}(\rho^{-5})\ .$ (97)
@@
The divergent behavior of $K$ in the standard case of Eq. (97) is also better than in the Reissner-Nordström geometry, since $K$ maintains the behavior $\rho^{-6}$ characteristic of the (chargeless) Schwarzschild geometry. Noticeably, this behavior can still be improved by choosing a mass such that  [^11] This mass is extremal for $\alpha=1$ (see Eq. (86)).
$C=\frac{2\ \Gamma(\frac{1}{4})\Gamma(\frac{5}{4})}{3\sqrt{\pi}\,\alpha}\ .$ (99)
@@
$K=\frac{4}{\epsilon^{2}\rho^{4}}-\frac{8}{3\epsilon^{2}\rho^{2}}+\frac{52}{15\,\epsilon^{2}}+\mathcal{O}(\rho^{2})\ .$ (101)
Instead, the behavior of $K$ in the anomalous case cannot be improved by choosing a particular value for $C$, since the first term in Eq. (96) does not depend on $C$. Neither the behavior of $K$ in Reissner-Nordström geometry could be improved in this way. Consistently, the same $r^{-4}$ dependence of the Kretschmann scalar was found in [38] solving the Einstein equations sourced by the BI electromagnetism, for a dyonic BH and a critical value of the mass.
The existence of a true singularity is confirmed by the fact that both $\rho=1$ in the anomalous case, or $\rho=0$ in the standard case can be reached in a finite proper time. This is because the behavior of $f$ at the singularity is either regular (anomalous case) or the same as in an ordinary BH (standard case). Therefore, spacetime is geodesically incomplete.
@@ VI Conclusions
Born–Infeld electrogravity is governed by a Lagrangian that couples gravity and electromagnetism through a unique determinantal structure, in which the Lagrangian density is given by the determinant of a tensor whose symmetric part is geometric and whose antisymmetric part is electromagnetic. The Palatini formalism extracts the dynamics by taking the metric $\bf g$, the affine connection $\bf\Gamma$, and the electromagnetic potential $\bf A$ as independent dynamical variables. It exhibits its full power by yielding well-behaved dynamic equations free of instabilities. The variation with respect to $\bf g$ leads to a constraint equation, which is the consequence of the absence of derivatives of the metric in the Lagrangian. In turn, the dynamical content of the theory comes from the variations with respect to $\bf\Gamma$ and $\bf A$. The variation with respect to $\bf\Gamma$ must preserve the symmetry attributed to the geometric sector; this is achieved by setting the torsion to zero. This condition, together with the constraint equation, leads to the metric compatibility and fixes the connection as the Levi-Civita connection.
@@
As they arise from the Palatini formalism, the field equations are separated in terms of the symmetric and antisymmetric parts of the auxiliary tensor $\bar{q}^{\mu\nu}$. The dynamics of $\bar{q}^{[\mu\nu]}$ is the standard Born–Infeld electrodynamics as developed over an effective geometric background defined by the symmetric tensor $\mathcal{G}_{\mu\nu}\equiv g_{\mu\nu}+\epsilon R_{(\mu\nu)}=q_{(\mu\nu)}$. Nevertheless, an electrodynamical picture in the background of the physical metric $g_{\mu\nu}$ is also realizable by combining the symmetric and antisymmetric sectors of the dynamics. This second description retains its Born–Infeld essence but exhibits a sign alteration in two terms that contain the quadratic invariants $S$ and $P$; therefore, we refer to it as “anomalous”. We have shown that these two pictures, $\mathcal{G}$” and “gg”, are related by a set of identities that map the pseudoscalars built with $\mathcal{G}$ into those built with $g$, and that also relate the corresponding volume elements (Eqs. (47), (49) and (51)). In particular, these relations explain how the apparent coupling of the electromagnetic field to curvature in picture $\mathcal{G}$ (Eq. (37)) is reabsorbed and translated into a minimally coupled Born–Infeld structure in picture $g$, in agreement with the equivalence principle. It is worth noting that the anomalous and standard characters of each picture are not essential since they can be interchanged by replacing $(\beta,\epsilon)$ with $(i\beta,-\epsilon)$, which is allowed because only even powers of $\beta$ appear in the Lagrangian. These two electrodynamical pictures, in the backgrounds $\mathcal{G}$ and $g$, resemble Plebański’s developments [24] that show that the same field configuration can be conceived in two different metrics related through the energy-momentum tensor of the field (see also [26]). In our case this is so, because $\epsilon R_{\mu\nu}$ in $\mathcal{G}_{\mu\nu}$ are connected to the electromagnetic energy-momentum tensor through the Einstein equations.
@@
To explore the implications of the model on the singular behavior of the solutions, we analyzed spherically symmetric dyonic configurations. If $\beta$ is real in the Lagrangian, then the electric field becomes singular at a finite radius $r_{o}$, restricting the domain of the solution to $r>r_{o}$, while the magnetic field maintains its Coulombian form. If instead the parameter $\beta$ is taken to be imaginary, the electric field acquires the standard Born–Infeld form but the magnetic monopole is divergent at the origin. Remarkably, even though the electric and magnetic monopoles enter the field invariants $S$ and $P$ unequally, they contribute to the energy and pressure on an equal footing through the combination $\sqrt{q^{2}+p^{2}}$, a feature that is inherited by the geometry. The obtained geometric solution approaches the Reissner–Nordström form asymptotically, in the weak gravity region. The family of solutions admits two, one (extremal) or no horizons depending on the mass and parameter choices, with thermodynamical properties typical of this type of BHs. The relation mass vs. charge for the extremal BH is similar to the one of the extremal Reissner–Nordström BH if the charge is large ($\alpha<1$) but differs significantly for small charge, because the mass of the BI extremal BH goes to $M_{\text{extr}}=\beta\,c^{4}/(4G^{3/2})$ for decreasing charge (increasing values of $\alpha>1$). In this limit, the mass and the horizon radious are characterized exclusively by the fundamental constants of the theory. The horizon radious of this fundamental extremal BH satisfies $r_{h}=2GM_{\text{extr}}$, like the Schwarzschild BH. However, unlike the Schwarzschild BH, its mass is not an integration constant but the natural unit of mass defined by the fundamental constants of the theory; besides, its temperature is zero, as is typical of an extremal BH.
@@
With regard to the singularities, Born–Infeld electrogravity alleviates the geometric divergences compared to the Reissner–Nordström case, but nevertheless spacetime remains geodesically incomplete. The improvement is more pronounced in the real-$\beta$ case, because the curvature singularity is shifted from the center to a sphere at $r=r_{o}$, mitigating the pathological inverse power behavior of the Kretschmann scalar. Notably, for imaginary $\beta$ there exists a solution for which the metric becomes regular at the origin and the singularity in the Kretschmann scalar is of order $r^{-4}$.
@@ Appendix B The projective mode
The antisymmetry in the last two indices of $R^{a}{}_{bcd}$ reflects the 2-form nature of $\widetilde{\widetilde{\mathcal{R}}}{}^{a}_{\;b}$, while the indices $a$, $b$ label the different members of the family. Let us now consider a projective transformation of the connection